A container is partly filled with 12 liters of whole milk, containing 4% butterfat, How much 1% milk must be added to get 2% butterfat
Answer by stanbon(75887)

A container is partly filled with 12 liters of whole milk, containing 4% butterfat, How much 1% milk must be added to get 2% butterfat
-----------------------------
Equation:
fat + fat = fat
0.04*12 + 0.01x = 0.02(12+x)
-----------
Multiply thru by 100 to get:
4*12 + x = 2*12 + 2x
x = 2*12
x = 24 liters (amt. of 1% milk needed for the mix)
==================================================
Cheers,
Stan H.
===============